OpenStudy (anonymous):

square root of x^3y^3 / square root xy square root of x^3y^3 / square root xy @Mathematics

OpenStudy (anonymous):

\[\sqrt{x^3y^3} / \sqrt{xy}\]

jhonyy9 (jhonyy9):

=sqrt(x2y2)=xy